Look at the numbers in the following list.
Identify the rational numbers
Negative square root of 16,Negative 3.14 (pie),Fraction 15 over 4,Fraction 14 over 3
Graph numbers on number line

Answer by midwood_trail(310)

Negative square root of 16,Negative 3.14 (pie),Fraction 15 over 4,Fraction 14 over 3
Graph numbers on number line.
-sqrt{16} = -4
-pi = -3.14
15/4 = 3.75
14/3 = 4.66 or 4.7
We now plot the numbers on the number line.

<---(-4)----(-3.14)---------(0)---------(3.75)--------(4.66)------>